OpenStudy (anonymous):

1. Evaluate the expression for the given values(s) of the variable(s) 8- 6x/y +2x for x=2 y=4 2.To rent a bicycle, pay a $12 fee plus $2 per hour. Write an algebraic expression to represent the total cost for renting a bike h hours a) 12+h/2 b)2h+12 c)14h d)2/h+12 3.Simplify the expression 54-6x3+4^2 4.Tell which property is represented. 17+(3+1)=(17+3)+1 a) commutative property b)associative property c)identity property d)distributive property 5.simplify 7(x+4)+2 a)7x+6 b)7x+30 c)14x+60 d)x+13 6. simplify the expression 4-24/2^3

OpenStudy (anonymous):

1) 8- 6x/y +2x for x=2 y=4 8-(6*2)/4+2*2= 8-12/4+4= 8-3+4=9 2) b)2h+12 3) 54-6x3+4^2= 54-18+16=20 4) a) commutative property 5) 7(x+4)+2= 7x+28+2= 7x+30 so the answer is b) 6) 4-24/2^3= 4-24/8= 4-3=1

OpenStudy (xxdragoonxx):